# Conformal Validity Guarantees Exist for Any Data Distribution

### v1

@@ 5.1 Multistep Protein Design Experiments
Efficient Computation with Split CP for Multistep Protein Design While the high training cost of Full CP is prohibitive for many types of ML predictors, this limitation is avoided by variants of Split CP, which we also extend to MFCS. Variants of Split CP reduce computational effort by splitting the available data into a “proper” training set that is only used to fit a single ML model, and a separate holdout “calibration” set that is evaluated with the fitted model (via the score function) to construct the CP set. Split CP’s computational complexity thus corresponds to the evaluation cost of the pre-fitted model on the calibration and test data.
@@ Appendix A Main Result Proof and Details
which follows since, conditioned on $E_{z}$, there are an equal number of scores $v_{1},...,v_{n+1}$ as data points $z_{1},...,z_{n+1}$, thus implying that the fitted (symmetric) score function $\mathcal{\widehat{S}}$ induces a bijection between the discrete set of distinct data observations and the discrete set of distinct scores. [^9] To be thorough and precise, the equation $\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{z}\}=\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$ relies on the convention in conformal prediction that $\mathcal{S}$ is a symmetric score function (see Remark 4.3). We elaborate on this point in Appendix A.3. Then, using Bayes rule and the law of total probability (LOTP), it follows that
@@
In Theorem 4.1 and the proof in Appendix A.1 we use score functions that correspond to the “ordinary” full conformal method Vovk et al. (2005). However, the same result and proof technique would also apply to generalizing the “deleted” or “leave-one-out” full conformal method (which is computationally more expensive than ordinary full CP but tends to obtain more informative prediction sets by avoiding overfitting in its construction) Vovk et al. (2005) or the split conformal method Papadopoulos (2008). The score functions for the deleted full CP method would be $V_{i}^{(x,y)}=\mathcal{S}(Z_{i},Z_{-i}\cup\{(x,y)\})$ for $i\in\{1,...,n\}$, where $Z_{-i}=Z_{1:n}\backslash Z_{i}$, and $V_{n+1}^{(x,y)}=\mathcal{S}\big{(}(x,y),Z_{1:n}\big{)}$. For split CP, the score functions could be defined using independent training and calibration sets. Letting $Z^{0}_{1:m}=\{Z_{i}^{0}\}_{i=1}^{m}$ denote the training data and $Z_{1:n}=\{Z_{i}\}_{i=1}^{n}$ denote the calibration data, the split CP score functions would be $V_{i}^{(x,y)}=\mathcal{S}(Z_{i},Z_{1:m}^{0})$ for $i\in\{1,...,n\}$ and $V_{n+1}^{(x,y)}=\mathcal{S}\big{(}(x,y),Z_{1:m}^{0}\big{)}$.
@@ Appendix B Deriving Weighted Conformal Validity Guarantees and Algorithms for Data under MFCS
$=\mathbb{P}(Z_{1}=z_{1},...,Z_{n+t}=z_{n+1})$ More explicit notation
@@
$\ \begin{matrix}[l]=\mathbb{P}(Y_{n+t}=y_{n+t}\ |\ Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1},X_{n+t}=x_{n+t})\\ \qquad\cdot\ \mathbb{P}(X_{n+t}=x_{n+t}\ |\ Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1})\\ \qquad\cdot\ \mathbb{P}(Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1})\end{matrix}$ Bayes or chain rule
@@ Appendix E Additional Experimental Results
E.1 Bounding the Query Function to Enforce Informative (Finite) Intervals
@@
MEPS dataset
Figure 6: Unbounded (top row) versus bounded (bottom row) active learning experiments of proposed multistep split CP method for $d=3$ (red circles) compared to baselines of unweighted split CP (orange squares), one-step split CP (green triangles), and ACI (gray squares) on the MEPS dataset. The Y-axes represent mean coverage, median interval width, and mean squared error on a holdout test set; the X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. All values are computed over 350 distinct random seeds.

### v2

@@ 5.1 Multistep Protein Design Experiments
Efficient Computation with Split CP for Multistep Protein Design While the training cost of Full CP is prohibitive for many types of ML predictors, this limitation is avoided by variants of Split CP (Papadopoulos, 2008), which we also extend to MFCS. Split CP methods reduce computational effort by splitting the available data into a “proper” training set used only to fit a single ML model, and a separate holdout “calibration” set on which the fitted model computes CP scores to construct the CP set. Split CP’s computational complexity thus corresponds to the evaluation cost of the pre-fitted model on the calibration and test data.
@@ Appendix A Main Result Proof and Details
which follows since, conditioned on $E_{z}$, there are an equal number of scores $v_{1},...,v_{n+1}$ as data points $z_{1},...,z_{n+1}$, thus implying that the fitted (symmetric) score function $\mathcal{\widehat{S}}$ induces a bijection between the discrete set of distinct data observations and the discrete set of distinct scores. [^10] To be thorough and precise, the equation $\mathbb{P}\{V_{n+1}=v_{i}\ |\ E_{z}\}=\mathbb{P}\{Z_{n+1}=z_{i}\ |\ E_{z}\}$ relies on the convention in conformal prediction that $\mathcal{S}$ is a symmetric score function (see Remark 4.3). We elaborate on this point in Appendix A.3. Then, using the definition of conditional probability and the law of total probability (LOTP), it follows that
@@
In Theorem 4.1 and the proof in Appendix A.1 we use score functions that correspond to the “ordinary” full conformal method Vovk et al. (2005). However, the same result and proof technique would also apply to generalizing the “deleted” or “leave-one-out” full conformal method (which is computationally more expensive than ordinary full CP but tends to obtain more informative prediction sets by avoiding overfitting in its construction) (Vovk et al., 2005) or the split conformal method (Papadopoulos, 2008). The score functions for the deleted full CP method would be $V_{i}^{(x,y)}=\mathcal{S}(Z_{i},Z_{-i}\cup\{(x,y)\})$ for $i\in\{1,...,n\}$, where $Z_{-i}=Z_{1:n}\backslash Z_{i}$, and $V_{n+1}^{(x,y)}=\mathcal{S}\big{(}(x,y),Z_{1:n}\big{)}$. For split CP, the score functions could be defined using independent training and calibration sets. Letting $Z^{0}_{1:m}=\{Z_{i}^{0}\}_{i=1}^{m}$ denote the training data and $Z_{1:n}=\{Z_{i}\}_{i=1}^{n}$ denote the calibration data, the split CP score functions would be $V_{i}^{(x,y)}=\mathcal{S}(Z_{i},Z_{1:m}^{0})$ for $i\in\{1,...,n\}$ and $V_{n+1}^{(x,y)}=\mathcal{S}\big{(}(x,y),Z_{1:m}^{0}\big{)}$.
@@ Appendix B Deriving Weighted Conformal Validity Guarantees and Algorithms for Data under MFCS
$=\mathbb{P}(Z_{1}=z_{1},...,Z_{n+t}=z_{n+t})$ More explicit notation
@@
$\ \begin{matrix}[l]=\mathbb{P}(Y_{n+t}=y_{n+t}\ |\ Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1},X_{n+t}=x_{n+t})\\ \qquad\cdot\ \mathbb{P}(X_{n+t}=x_{n+t}\ |\ Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1})\\ \qquad\cdot\ \mathbb{P}(Z_{1}=z_{1},...,Z_{n+t-1}=z_{n+t-1})\end{matrix}$ Chain rule or conditional probability definition
@@ Appendix E Additional Experimental Results
E.1 Bounding the Query Function to Enforce Informative (Finite) Prediction Intervals
@@
Medical Expenditure Panel Survey (MEPS) dataset
Figure 6: Unbounded (top row) versus bounded (bottom row) active learning experiments with proposed multistep split CP method for estimation depth $d=3$ (red circles) compared to baselines of unweighted split CP (orange squares), one-step split CP (green triangles), and ACI (gray squares) on the Medical Expenditure Panel Survey (MEPS) dataset. The Y-axes represent mean coverage, median interval width, mean squared error on a holdout test set, and the relative bound magnitude $B/B_{\text{max}}$; the X-axes correspond to the number of active learning query steps, with each query based on posterior variance of a GP regressor. All values are computed over 350 distinct random seeds.